[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Apr 18, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Apr 18 14:55:04 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig

----------

TAFB:
1. TSB Developer Matt Sardi investigatied the problem of missing radar  
imagery in N-AWIPS and discovered multiple symbolic links and disk mounts  
(in /etc/auto.direct) missing on 5 operational Linux workstations in TAFB  
(cumulus, starfish, ballyho, shiner, grunt) and 5 operational workstations  
in HSU (sawfish, ridges, tilefish, hurrback, bonefish). He submitted an  
URGENT NCO helpdesk ticket to get them reestablished. NCO Linux Sysadmin  
Brian Maher fixed the problems and restored access to the data. TAFB  
forecasters report that all data is now available in N-AWIPS.

2. TAFB forecasters at the Surface Analysis desk are continuing to use  
Sheephead and New Crappie until the Cumulus workstation is stablilized by  
NCO after the RHEL 6 upgrade.
HSU:
1. A P-Surge test was run for fake storm AL802017 at 1030 am EDT (1430Z)  
today.
Other items:
1. The cause of yesterday's network outage was a switch upgrade (network  
traffic moved from router NCOCR1 to NCOCR2) in College Park's Data Center  
and a subsequent failure of multiple DNS servers as part of RFC #3039.  
Dataflow was interrupted at many national centers as well as Environment  
Canada. Original estimate of impact was 30-60 seconds; work on lingering  
problems continues this morning.
NHC RFCs:
1. RFC #3189 - Make changes to IP tables on the NHC systems on the 112  
VLAN. An NHC user has requested SSH Access from the nhcgis.ncep.noaa.gov  
and nhcgisdev.ncep.noaa.gov systems to the 176 VLAN. To be implemented on  
Tuesday, April 18 at 1600Z (noon EDT).

2. RFC #3192 - Make a change to allow user atcf-ops to transfer files from  
NHC's new ATCF server (atcfsvr3) to the Opah data server via SSH. To be  
implemented on Wednesday, April 19 at 1400Z (10 am EDT).

3. RFC #3153 - Move the NCEP network traffic from the global routing table  
to the NCEP VRF (virtual routing and forwarding) instance. This is the last  
step to allow NCO to route oneNWSnet (WFO) traffic between Boulder/College  
Park to Reston/Orlando and exchange routes with existing networks (legacy  
opsnet). Originally scheduled for Tuesday, April 11 at 1100Z (7 am EDT),  
postponed indefinitely (TBD).

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170418/24a1f2fe/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list